# Chaos Engineering: A Multi-Vocal Literature Review

### v1

@@ 3.2.4. Security and Governance
Starting with known knowns, teams establish a baseline of predictable system behavior. They can then progressively move into more uncertain areas like Known Unknowns and Unknowns Knowns before finally addressing the most unpredictable Unknown Unknowns (Palani and Gupta, 2023). This step-by-step approach reduces system risks by building understanding before testing less predictable areas.
Ensure Regulatory Compliance.
Adhering to industry standards and legal frameworks is essential to ensure that systems meet established data protection and security requirements, including guidelines such as the General Data Protection Regulation (GDPR) and IEC 62443 (Helmke, 2019; Torkura et al., 2021). These frameworks provide organizations with critical guidance to safeguard sensitive information, maintain operational integrity, and protect vital infrastructures (Helmke, 2019; Kostic, 2024). To meet these compliance requirements, some organizations adopt Infrastructure as Code (IaC), which automates the provisioning and management of infrastructure (Torkura et al., 2020; Fogli et al., 2023). Tools such as Terraform and Ansible are commonly used in IaC to ensure consistent implementation of security policies, minimize human error, and maintain infrastructure in alignment with regulatory standards (Torkura et al., 2021; Vu et al., 2022). In addition to IaC, monitoring tools such as Datadog  (Abdul, 2024; Mooney, 2023), which detects compliance violations and configuration drift such as publicly exposed data, and Security Monkey, which monitors Secure Sockets Layer (SSL) certificates and other security settings to prevent vulnerabilities such as certificate expiration, are essential to maintain compliance (Dua, 2024; Tivey et al., 2019).
3.2.5. Integration and Automation with Continuous Integration and Continuous Delivery (CI/CD)
CI/CD pipelines are critical to optimizing the efficiency and consistency of chaos experiments (Rivera et al., 2023; Bairyev, 2023; Azure Readiness, 2023). When performed manually, tasks such as performing experiments and monitoring system behavior can be inefficient and prone to human error (Kostic, 2024; Jernberg et al., 2020; Bremmers, 2021; Chen et al., 2024). Automating these processes minimizes human intervention, ensuring tasks are performed reliably and consistently, thus reducing the likelihood of mistakes (Wickramasinghe, 2023; Camacho et al., 2022; Simonsson et al., 2021; Green, 2023). Integrating chaos experiments into CI/CD pipelines further enhances this reliability by automatically initiating tests as part of the routine development workflow (Roa, 2022). With each new deployment or configuration update, chaos tests run alongside other routine tests, enabling early detection of vulnerabilities. This proactive approach helps ensure issues are addressed in the development phase, resulting in a more comprehensive and practical testing process (Bremmers, 2021; Manna, 2021)
Figure 5. Key Functionalities Provided by Chaos Engineering.
[…]
Functional Requirement
3.3. Quality Requirements for Chaos Engineering Platform
We have identified 12 quality requirements for a chaos engineering platform from the selected literature: performance, recoverability, resilience, timeliness, observability, scalability, effectiveness, robustness, accuracy, fault tolerance, cost, and adaptability. We provide these requirements and the corresponding metrics for measuring them in Table 8.
Table 8. Quality Attributes and Measurement Metrics
Quality Attribute Measurement Metrics Articles Referenced Total
Performance Evaluate resource use and performance under attack, ensuring chaos experiments align with production metrics. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Nikolaidis et al., 2021; Vu et al., 2022; Klinaku et al., 2022) 10
Recoverability Measure mean time to repair (MTTR), restoration speed, and detection response. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Camacho et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Zhang et al., 2023; Tomka, 2024; Kadikar, 2023) 16
Resilience Tolerance for degraded performance under stress; success rate of automated restoration actions. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Nikolaidis et al., 2023; Frank et al., 2023; Kassab et al., 2022; Konstantinou et al., 2021; Nombela, 2023; Krishnamurthy, 2021; Bocetta, 2019; Manna, 2021; Kostic, 2024; Krivas and Portela, 2020; National Australia Bank, 2020; Azure Readiness, 2023) 20
Timeliness Time to restore service and detect interruptions; impact on user experience and disruption. (Camacho et al., 2022; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Cotroneo et al., 2022; Basiri et al., 2019; Gill, 2021; Zhang et al., 2019; Durai, 2022; Shah, 2021; Katirtzis, 2022; Dua, 2024; Wu, 2021; Palani and Gupta, 2023; Manna, 2021) 16
Observability Assess logging completeness and monitoring uptime; measure alert latency and tracing success rate. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Frank et al., 2023; Malik et al., 2023; Kesim et al., 2020; Siwach et al., 2022; Konstantinou et al., 2021; Hawkins, 2020; Manna, 2021; Kostic, 2024; Kadikar, 2023; Azure Readiness, 2023) 20
Scalability Manage increasing attack rates and scale experiment complexity. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Patel, 2022; Gill, 2021; Vu et al., 2022; Sondhi et al., 2021; Malik et al., 2023; Klinaku et al., 2022; Camilli et al., 2022; Nikolaidis et al., 2023; Kassab et al., 2022; Rivera et al., 2023; Krishnamurthy, 2021; Durai, 2022; Wu, 2021; Palani and Gupta, 2023; Manna, 2021; Azure Readiness, 2023; Vinisky, 2024; Kostic, 2024) 20
Effectiveness Measure achievement rates, end-user satisfaction, and time to achieve outcomes. (Sondhi et al., 2021; Chen et al., 2022; Malik et al., 2023; Wu, 2021) 4
Robustness Evaluate stability at peak load and error tolerance. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021; Bedoya et al., 2023; Chen et al., 2022; Vinisky, 2024) 7
Accuracy Evaluate system consistency through recovery accuracy and fault localization precision. (Green, 2023; Simonsson et al., 2021; Chen et al., 2022) 3
Fault Tolerance Automated failover success rate during faults; immediate recovery time after detected faults. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021a; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Ma et al., 2023; Meiklejohn et al., 2021; Sondhi et al., 2021; Azure Readiness, 2023; Tomka, 2024; Manna, 2021) 15
Cost Cost of chaos engineering and savings from preventing outages. (Gill, 2021; Patel, 2022; Kesim et al., 2020; Palani and Gupta, 2023; Klinaku et al., 2022; Inc., 2023; Bocetta, 2019; Basiri et al., 2019; National Australia Bank, 2020) 9
Adaptability Change methods and platforms for new systems. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021) 4
3.4. Key Components of Chaos Engineering
Before adopting a chaos engineering framework, organizations should carefully evaluate its key components to ensure that they align with their operational requirements and objectives. We have identified several essential components for an effective chaos engineering platform. Most of the reviewed studies incorporate a combination of these components, which we have categorized into five central units: experimentation, collaboration and communication, observability and analysis, security and governance, and automation and integration with CI/CD. This classification, derived from the functionalities discussed in Section  3.2, is shown in Figure  6. The remainder of this section provides a detailed analysis of each component within this framework.
@@ 3.4. Key Components of Chaos Engineering
3.5. Experimentation Unit
Successful coordination of chaos experiments depends on a systematic approach to both planning and execution (Zhang et al., 2021a; Nikolaidis et al., 2021). To this end, the experimentation unit is built around two foundational components: the experiment designer and the failure injection hub, which will be explored further in this section.
3.5.1. Experiment Designer
This module is crucial for structuring chaos experiments, allowing teams to define which systems to target (Zhang et al., 2021a; Camacho et al., 2022). After identifying the systems, the teams can choose the fault types most relevant to their testing scenarios, such as network failures or system crashes (Zhang et al., 2021a). Once the faults are defined, the experiment designer helps organize the timing and order in which they are introduced, ensuring that they reflect real-world conditions (Camacho et al., 2022). By enabling these customized failure simulations, the experiment designer ensures that the tests are well-structured, providing consistent data for postmortem analysis (Zhang et al., 2021a; Konstantinou et al., 2021). This structure allows teams to track how the behavior of the system evolves under different failure conditions. For example, teams might simulate a server crash during peak traffic to test failover capabilities and measure the impact on user experience (Zhang et al., 2019; Simonsson et al., 2021).
3.5.2. Fault Injection Hub
Once the experiment is designed and the faults are organized, the next phase is to inject these controlled faults and workloads into the system, simulating real-world scenarios (Fawcett, 2020; Palani and Gupta, 2023). The fault injection process, as outlined by Hseh et al (Hsueh et al., 1997), in Figure  7, involves multiple components working together: a controller, fault injector, fault library, workload generator, workload library, and target system (Nikolaidis et al., 2021; Frank et al., 2021b). Each of these components plays a critical role in ensuring the accuracy and smooth execution of the chaos experiment. Central to this process is the controller, which manages the introduction of faults and dynamically adjusts based on real-time system feedback, ensuring that the experiment remains aligned with its objectives (Torkura et al., 2020). As part of the hub, the fault injector introduces specific disruptions, such as hardware or software failures, selected from a diverse fault library that offers teams a variety of failure scenarios to choose from (Soldani and Brogi, 2021; Frank et al., 2021b). In parallel, the workload generator creates synthetic workloads based on predefined patterns stored in the workload library. This allows teams to simulate normal or peak traffic conditions to observe how the system handles varying stress levels (Fogli et al., 2023; Kesim et al., 2020). These faults and workloads are then applied to the target system, be it a production or preproduction environment, providing teams with valuable information on the weaknesses of the system and the recovery capabilities (Wickramasinghe, 2023; Zhang et al., 2021a).
Figure 7. Conceptual Schema of Fault Injection(Hsueh et al., 1997).

### v2

@@ 4.2.1. Experiment Design and Planning
. This framework supports a gradual testing strategy—starting with predictable behaviors and extending to more uncertain or unforeseen failure modes. Doing so helps reduce risk while expanding organizational understanding of system weaknesses (Palani and Gupta, 2023a; Inc., 2023; Chen et al., 2022).
4.2.2. Fault Execution and Control
This phase ensures that faults are introduced in a controlled, safe, and purposeful manner to validate the system’s resilience under adverse conditions. The goal is to simulate real-world failures while minimizing risk and maximizing learning (Zhang et al., 2021b; Green, 2023; Patel, 2022).
Inject Faults Based on Defined Scenarios
Chaos experiments begin by executing predefined fault scenarios aligned with the experiment plan. These include disruptions such as service crashes, network partitions, and CPU saturation (Zhang et al., 2021b; Cotroneo et al., 2022; Nikolaidis et al., 2023; Siwach et al., 2022). Faults are introduced progressively—starting with a minimal scope and gradually increasing impact—to monitor system behavior and avoid widespread disruption (Fogli et al., 2023; Wickramasinghe, 2023; Simonsson et al., 2021; Chen et al., 2024; Green, 2023). Execution readiness checks ensure that failure types are verified, system health is stable, and rollback plans are in place (Zhang et al., 2019; Vu et al., 2022; Chen et al., 2022; Kadikar, 2023; Bairyev, 2023).
Apply Safety Mechanisms
To protect against unintended system degradation, safety measures are implemented. These include automatic abort thresholds that halt experiments if system performance degrades beyond acceptable levels (Frank et al., 2023; Malik et al., 2023; Chen et al., 2024; Camacho et al., 2022). Rollback protocols ensure the system can return to a known steady state, while blast radius limits prevent faults from affecting critical components (Simonsson et al., 2021; Green, 2023; Aharon, 2024; Inc., 2023; Kadikar, 2023).
Notify Stakeholders During Experiment Lifecycle
Transparent communication is maintained throughout the experiment. Key stakeholders are informed at defined checkpoints (start, progress, and end of experiment) to facilitate oversight, enable timely intervention, and foster cross-functional trust in the experimentation process (Torkura et al., 2020; Siwach et al., 2022; Patel, 2022; Bairyev, 2023; Le, 2022; Green, 2023; Zhang et al., 2019).
4.2.3. Monitoring and Observability
This phase enables teams to understand how the system behaves under stress by continuously collecting and interpreting telemetry data during chaos experiments. Its purpose is to detect deviations, validate resilience assumptions, and guide response strategies (Zhang et al., 2019; Camacho et al., 2022; Inc., 2023; Nikolaidis et al., 2021).
Capture Live System Telemetry
The process begins with configuring telemetry tools to collect real-time data, such as metrics, logs, and traces (Simonsson et al., 2021; Pulcinelli et al., 2023; Nikolaidis et al., 2021). These data streams encompass both application-level indicators (e.g., service availability, request latency) and infrastructure metrics (e.g., CPU usage, memory consumption) (Gill, 2021; Chen et al., 2024). To ensure full system visibility, team members are assigned specific monitoring scopes—some focusing on business-critical services, others on backend components or network behavior (Vu et al., 2022; Nikolaidis et al., 2023). When unexpected behaviors occur, teams follow predefined diagnostic steps, such as analyzing error logs, correlating metrics, and reviewing recent deployment changes, to progressively isolate and identify the faulty components (Ma et al., 2023; Sondhi et al., 2021; Patel, 2022; Chen et al., 2022).
Visualize Experiment Outcomes and Behaviors
Collected data is visualized through dashboards that track key metrics over time and flag deviations from expected behavior (Inc., 2023; Gill, 2021; Nikolaidis et al., 2021; Chen et al., 2022; Bairyev, 2023; Poltronieri et al., 2022). These visual tools guide real-time decisions and help verify whether system behavior aligns with predefined hypotheses (Zhang et al., 2019; Butow, 2018; Bremmers, 2021; Nikolaidis et al., 2023). Observations are documented with relevant metadata, such as timestamps, environment context, and test conditions—to support future analysis and continuous learning (Naqvi et al., 2022; Sondhi et al., 2021; Dedousis et al., 2023; Green, 2023)
4.2.4. Post-Experiment Analysis
This phase focuses on deriving actionable insights from chaos experiments by assessing impact, diagnosing root causes, and preserving knowledge for future use (Chen et al., 2022; Bremmers, 2021; Frank et al., 2023; Patel, 2022; Al-Said Ahmad et al., 2024). The process involves three key areas of functionality.
Analyze Impact and Diagnose Root Causes
After fault injection, teams analyze telemetry, such as metrics, logs, and traces, to compare system performance against baseline expectations (Simonsson et al., 2021; Vu et al., 2022; Inc., 2023; Dedousis et al., 2023). Deviations are measured to understand the extent of impact on key indicators such as latency or availability (Frank et al., 2023; Poltronieri et al., 2022; Jernberg et al., 2020). This helps prioritize which failures matter most. Root cause analysis then investigates how the failure unfolded, using traces and service maps to locate the origin and propagation path of faults (Chen et al., 2024; Ma et al., 2023; Zhang et al., 2021a, 2019; Bremmers, 2021; Bairyev, 2023).
Archive and Document Learnings
All findings are documented in structured postmortem reports, which capture the fault scenario, affected services, violated hypotheses, and recovery behavior (Vu et al., 2022; Kadikar, 2023; Bairyev, 2023; Jernberg et al., 2020; Chen et al., 2022). These insights are stored in internal knowledge bases to support incident response and guide future experiments (Zhang et al., 2019; Siwach et al., 2022; Green, 2023). By archiving this information, teams enable iterative learning and systemic improvements in both technical and organizational resilience (Green, 2023; Dedousis et al., 2023).
4.2.5. Automation and Integration
This phase automates chaos experiments to minimize manual effort and then integrates them into delivery pipelines and platform workflows, ensuring that resilience testing becomes routine.
Automate Chaos Testing Pipelines
Automation reduces manual effort by enabling chaos experiments to run consistently across different stages of the software delivery process, such as development, staging, and production (Chen et al., 2022; Pierce et al., 2021; Bedoya et al., 2023). Teams can define test parameters and execution policies in advance, ensuring that faults are injected at predictable times or in response to deployment events (Manna, 2021; Bairyev, 2023). This supports repeatability, reduces errors, and allows for continuous resilience validation throughout the development lifecycle (Gill, 2021; Ikeuchi et al., 2020).
Integrate with platform tools
To embed chaos engineering into day-to-day engineering workflows, tools are integrated with existing platforms such as CI/CD (Continuous Integration and Continuous Delivery/Deployment) pipelines (e.g., Jenkins, GitHub Actions), cloud environments (e.g., Kubernetes, AWS), and observability frameworks (e.g., Prometheus, Grafana) (Simonsson et al., 2021; Manna, 2021; Camacho et al., 2022; Park et al., 2019). This integration enables the automated execution of chaos experiments during the build, test, and deployment phases, while also ensuring that experiment telemetry is collected and visualized in real-time. As a result, teams can detect weaknesses early in the lifecycle and improve response times during failure scenarios (Gill, 2021; Siwach et al., 2022; Bairyev, 2023; Nikolaidis et al., 2021).
4.3. Key Components of Chaos Engineering (RQ1.2)
Before adopting a chaos engineering platform, organizations should carefully evaluate its key components to ensure that they align with the operational requirements and objectives of the organization. Based on our systematic thematic analysis, we identified five primary components: the Experiment Design Unit, Fault Injection Unit, Observability Unit, Post-Experiment Analysis Unit, and Automation and Integration Unit (see Figure 5). This classification is based on the functionalities discussed in Section 4.2, mapping each component to the corresponding phase of the experimentation lifecycle. Each unit encapsulates modular services, such as hypothesis management, failure orchestration, observability pipelines, and automation engines, that together support scalable, safe, and repeatable experimentation. The remainder of this section details each component, its sub-modules, and their roles within the broader platform architecture. For details on coding, theme development, and mapping, see our online appendix (see Section 3.6).
@@ 4.2.5. Automation and Integration
4.3.1. Experiment Design Unit
This architectural component transforms high-level test objectives into system-executable configurations (Soldani and Brogi, 2021; Zhang et al., 2021b). It offers formal interfaces to define hypotheses, select test targets, and validate readiness prior to chaos injection (Sbiai et al., 2023; Chen et al., 2024; Dedousis et al., 2023; Frank et al., 2021b; Azure Readiness, 2023). The experiment design unit comprises two integrated modules: the Hypothesis Manager and the Target Selector.
Hypothesis Manager.
This enables teams to define, store, and manage test hypotheses through structured templates that capture expected system behavior using time-series metrics (e.g., latency, availability) and other indicators such as error rates and log patterns (Gianchandani et al., 2022; Bhaskar, 2022; Kadikar, 2023; Green, 2023; Le, 2022). Each hypothesis is programmatically linked to safety controls, including abort thresholds and rollback triggers, via policy configurations (Chen et al., 2024; Simonsson et al., 2021; Poltronieri et al., 2022). This ensures automated enforcement of safety during fault injection and accurate post-experiment validation (Zhang et al., 2021b; Dedousis et al., 2023).
Target Selector.
This phase scopes where faults will be injected by allowing teams to define experiment boundaries using filters such as environment (e.g., staging, production), infrastructure layer (e.g., compute, network), and service tags or labels (Soldani and Brogi, 2021; Azure Readiness, 2023; Green, 2023). It ensures controlled impact by enforcing blast radius limits, so only approved targets are affected (Sbiai et al., 2023).To maintain operational safety, it performs validation checks on all configurations, verifying correctness, authorization, and compatibility before execution. This prevents unsafe or unintended experiments from proceeding (Zhang et al., 2021b; Cotroneo et al., 2022; Bairyev, 2023; Chen et al., 2024).
4.3.2. Fault Injection Unit
This is the core architectural component responsible for executing chaos experiments by introducing faults into the system under test. It is composed of three core modules: the Fault Execution Engine, the Safety Controller, and the Alert Manager, each supporting a distinct aspect of the fault injection lifecycle.
Fault Execution Engine
This component orchestrates the actual injection of predefined faults and workloads. Its design follows the foundational model by Hsueh et al. (Hsueh et al., 1997) (See Figure 6), and has been extended to suit modern, distributed environments (Soldani and Brogi, 2021; Frank et al., 2021b). It comprises several modules: (1) A Controller that sequences experiment steps and responds to telemetry, (2) A Fault Injector that applies disruptions from a configurable fault library, (3) A Workload Generator that simulates system activity using patterns from a workload library, (4) The Target System, which receives the injected faults and workloads and emits telemetry for analysis (Wickramasinghe, 2023). Other components from the original model (see Figure 6)—such as the Monitor—are represented in the subsequent architectural component through the Observability Unit, which is responsible for telemetry collection. Similarly, the Data Collector and Chaos Analyzer are addressed within the Post-Experiment Analysis Unit, which supports impact assessment and root cause diagnosis following fault injection.
Safety Controller
This enforces runtime protection by embedding safeguards directly into the fault injection lifecycle (Zhang et al., 2021b; Cotroneo et al., 2022; Bhaskar, 2022; Lenka et al., 2018). It comprises three subsystems:
  • Blast Radius Control. This limits the impact of injected faults by using traffic routing and isolation rules. For instance, in Kubernetes, these rules restrict disruptions to specific services, namespaces, or pods within the system (Soldani and Brogi, 2021; Wu, 2021; Vu et al., 2022; Kesim et al., 2020).
  • Abort Monitor. This continuously checks key metrics (e.g., error rate, response time) and halts the experiment if thresholds are breached (Kadikar, 2023; Green, 2023; Le, 2022).
  • Rollback Handler. This restores the system to a steady state by triggering recovery workflows or replaying system snapshots (Cotroneo et al., 2022; Soldani and Brogi, 2021; Kadikar, 2023; Bhaskar, 2022).
Alert Manager
This component coordinates communication with stakeholders by triggering real-time notifications across channels (e.g., Slack, email) at critical points: start, escalation, and end (Gill, 2021; Green, 2023). It also enforces escalation policies when thresholds are breached to ensure rapid response (Patel, 2022; Azure Readiness, 2023; Camacho et al., 2022; Roa, 2022).
Figure 6. Conceptual Schema of Fault Execution Engine(Hsueh et al., 1997).